The demand and supply functions of a product are given by the following

equations: Qd =1500 - 2P and Qs = -1300 + 2P

The government announces a program to support a price increase of $30 per kg of

this grain, which imposes a price floor of $730.

A. What are the equilibrium price and quantity before price support policy?

B. What quantity of grain is purchased by the consumers, supplied by the

producers and purchased by the government at the support price?

C. What is the change in consumer surplus, producer surplus and total surplus?

What is the cost of government to implement this price support policy?

D. Later on, the government changes the price support policy and provides

subsidy of $20 per kg sold. What is the price paid by buyers, price received by

sellers, change in consumer surplus, change in producer surplus and

government cost?


1
Expert's answer
2022-04-11T12:10:14-0400

"Q_d=1500-2p,"

"Q_s=-1300+2p,"

"Q_d=Q_s,"

"1500-2p=-1300+2p,"

"2800=4p,"

"p=700,"

"Q_d=Q_s=100."